OpenStudy (consliver19):

Integrals Suppose f″(x)=−36sin(6x) and f′(0)=−3, and f(0)=−6. Find, f(π/4)=

OpenStudy (consliver19):

I guess what I'm asking is, do I take the antiderivative, and plug in the value of the previous derivative? I dont know if this makes sense.. sorry,

OpenStudy (danjs):

yes, recall when you take an integral how it has the +C constant that is included?

OpenStudy (danjs):

Those given points are the conditions to figure out what the constant will be.

OpenStudy (danjs):

\[\large f'(x)= \int\limits -36\sin(6x)dx = 6 \cos(6x)+C\] That is a family of curves for f'(X) depending on what C is. They give you a point f'(0)=-3 that is on the f'(x) f ' (x) = 6cos(6x) + C -3 = 6cos(6*0) + C C=-9 so f ' (x) = 6xos(6x) - 9

OpenStudy (danjs):

same thing to get the function y

OpenStudy (consliver19):

Yo, sorry for the super late response but thanks! Does the # term derivative matter at all?

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)